# AnyEdit: Edit Any Knowledge Encoded in Language Models

### v2

@@ Abstract
Large language models (LLMs) often produce incorrect or outdated information, necessitating efficient and precise knowledge updates. Current model editing methods, however, struggle with long-form knowledge in diverse formats, such as poetry, code snippets, and mathematical derivations. These limitations arise from their reliance on editing a single token’s hidden state, a limitation we term “efficacy barrier”. To solve this, we propose AnyEdit, a new autoregressive editing paradigm. It decomposes long-form knowledge into sequential chunks and iteratively edits the key token in each chunk, ensuring consistent and accurate outputs. Theoretically, we ground AnyEdit in the Chain Rule of Mutual Information, showing its ability to update any knowledge within LLMs. Empirically, it outperforms strong baselines by 21.5% on benchmarks including UnKEBench, AKEW, and our new EditEverything dataset for long-form diverse-formatted knowledge. Additionally, AnyEdit serves as a plug-and-play framework, enabling current editing methods to update knowledge with arbitrary length and format, significantly advancing the scope and practicality of LLM knowledge editing.
@@ 1 Introduction
Figure 1: Comparison of current methods and our AnyEdit. (a) and (d) illustrate the editing processes; (c) and (e) show the editing efficacy as the number of tokens within the to-be-updated knowledge increases; (b) and (f) depict the type of knowledge that each method can edit. Best viewed in color.
@@
Despite their success, existing model editing methods mostly face significant limitations in the length and diversity of the knowledge they can update. As shown in Figure 1 (c), even leading methods like AlphaEdit (Fang et al., 2024) and RECT (Gu et al., 2024) struggle to handle updates exceeding 100 tokens. Worse still, most methods are restricted to knowledge represented as structured (subject, relation, object) triples. However, real-world knowledge is often encoded in diverse formats (e.g., mathematical derivations and code snippets as shown in Figure 1 (f)) and frequently exceeds the 100-token threshold (Wu et al., 2024). These constraints are ill-suited for real-world scenarios, significantly narrowing the scope of model editing and hindering its broader advancement.
@@
Here we first conduct an in-depth analysis to identify why current methods fail for long-form diverse-formatted knowledge. Considering Figure 1 (a) again, existing methods typically rely on locating a single token, assuming that altering its hidden states will suffice to edit the LLM’s output (i.e., enabling the model to generate desired outputs reflecting new knowledge). However, long-form diverse-formatted knowledge is inherently more complex and information-dense than a single triplet, often requiring the integration of multiple critical tokens and intricate interdependencies among their hidden states. Thus, altering just a single token’s hidden state is often insufficient to ensure consistent and accurate knowledge generation. We formalize this limitation as the efficacy barrier of single-token editing [^1] The “Efficacy” is the metric proposed by the model editing method ROME (Meng et al., 2022), aiming to evaluate the success rate of eding. For more details, please refer to Appendix A.2., which is empirically validated in Section 3. As a consequence, existing methods remain constrained by the paradigm of locating a single token and, as a result, are unable to overcome the aforementioned limitation.
@@
Hence, a critical question naturally arises for updating long-form diverse-formatted knowledge: “Can multiple tokens be jointly located and edited to enable coherent knowledge updates?” A straightforward solution is to directly extend single-token editing to multiple tokens, however, it risks interference or conflicts between hidden state perturbations, undermining the coherence of to-be-updated knowledge and causing the performance drop as Figure 1 (c) showcases. In sight of this, we introduce AnyEdit, an autoregressive editing paradigm that enables collaborative token-level updates. As illustrated in Figure 1 (d), AnyEdit decomposes long-form knowledge into sequential chunks, treating each as independent sub-knowledge. During editing, we iteratively (1) locate the final token of the current chunk and (2) perturb its hidden states to maximize the likelihood of generating the subsequent chunk. Building on the Chain Rule of mutual information (Dobrushin, 1963) in Information Theory (Kullback, 1997), we theoretically demonstrate that this autoregressive process ensures the generation of consistent, complete long-form knowledge. AnyEdit offers two key advantages: (1) Adaptivity: The number of edited tokens can be adaptively adjusted with the knowledge length, avoiding redundant edits. (2) Generality: It supports diverse knowledge formats (e.g., poetry, code, math) by decoupling structure-specific constraints. By shifting from single-token to multi-token collaborative editing, AnyEdit overcomes the efficacy barrier of current methods, demonstrating the potential for practical, precise updates across diverse knowledge formats.
@@
To validate AnyEdit, we conduct a comprehensive evaluation against leading model editing methods (e.g., MEMIT (Meng et al., 2023), AlphaEdit (Fang et al., 2024), and UnKE (Deng et al., 2024)) on the prevailing LLMs such as Llama3-8B-Instruct [^2] https://llama.meta.com/llama3 and Qwen2.5-7B-Chat (Yang et al., 2024). Beyond standard benchmark datasets (e.g., Counterfact (Meng et al., 2022) and ZsRE (Meng et al., 2023)) that represent knowledge as triples, we curate EditEverything, a new benchmark for long-form knowledge in diverse formats. As shown in Figure 1 (f), this dataset includes entries up to 458 tokens — over twice the length of the longest sequences in existing benchmarks (e.g., 156 tokens in AKEW (Wu et al., 2024)) — and spans multiple domains, including mathematics, news, code, and biochemistry. Results on EditEverything and standard benchmarks demonstrate that AnyEdit surpasses all baselines, achieving a 21.5% average improvement in editing accuracy with comparable computational overhead. Furthermore, as the first autoregressive editing framework, AnyEdit also enables seamless integration with existing locate-then-edit methods, as shown in Figure 1 (e). This plug-and-play capability equips traditional approaches with the ability to handle knowledge with arbitrary length and format, significantly broadening the scope and practicality of LLM knowledge editing.
@@
Figure 2: Relationship between knowledge format, original output probability, and efficacy when applying advanced editing methods to update triplet-structured and diverse-formatted knowledge. For each category, we randomly sample 200 knowledge instances to conduct experiments. Best viewed in color.
@@
Figure 3: Relationship between the number of tokens in to-be-updated knowledge, probability shift under random input perturbations, and editing efficacy. We conduct experiments by truncating the sampled knowledge instances to enable editing across different token lengths. The lighter-colored bands represent variance. Best viewed in color.
@@ 2 Preliminary
where $\mathbb{P}_{f({\bm{h}}_{t}+\bm{\hat{\delta}})}$ represents the output probability when replacing ${\bm{h}}_{t}$ with ${\bm{h}}_{t}+\bm{\hat{\delta}}$ in the LLM. Finally, the LLM parameters are updated such that, given the input $X$, the hidden state of the key token is aligned with ${\bm{h}}_{t}+\bm{\delta}$. For more details, please refer to Appendix B.1.
3 Exploring Limitations of Existing Paradigm
Although single-token editing methods have been extensively studied in recent years (Fang et al., 2024; Wang et al., 2024; Deng et al., 2024; Gu et al., 2024; Wu et al., 2024), we argue that they are fundamentally limited in updating long-form, diverse formatted knowledge. In this section, we analyze and empirically validate these two limitations.
3.1 Limitation of Editing Diverse-formatted Knowledge
@@
3.2 Limitation of Editing Long-form Knowledge
Recent studies suggest that although LLMs leverage attention mechanism (Wang & Komatsuzaki, 2021), the dependency between distant tokens weakens as their positional distance increases (Liu et al., 2024). As a result, for long outputs $Y$ (e.g., $Y$ exceeding 100 tokens), perturbations applied to input tokens may have a diminishing influence on the later tokens within $Y$. In such cases, the shift of generation probability of $Y$ introduced by the perturbation $\bm{\delta}$ may be too small, making it insufficient to increase the probability of $Y$ above that of any other potential output.
To validate this, we repeat the experiments described in Section 3.1. Additionally, we use causal tracing (Meng et al., 2022), a common strategy in model editing, to quantify the shift in the generation probability of $Y$ introduced by the perturbation on $X$ (details of causal tracing can be found in Appendix B.1). Figure 3 illustrates the relationship between: the number of tokens in $Y$, the probability shift, and editing efficacy. The results demonstrate that long-form knowledge, which is less affected by single-token edits on the input, shows poor editing efficacy. In other words, the low probability shift introduced by single-token edits emerges as the inherent limitation in effectively updating long-form knowledge.
@@ 5 Experiments
Llama3-8B-It Pre-edited 63.18$\pm$0.38 23.67$\pm$0.52 62.73$\pm$0.31 23.52$\pm$0.51 64.03$\pm$0.32 15.74$\pm$0.42 40.20$\pm$0.46 5.52$\pm$0.10 65.77$\pm$0.37 16.25$\pm$0.47
FT-L 40.31$\pm$0.45 11.39$\pm$0.56 37.29$\pm$0.41 8.51$\pm$0.51 42.89$\pm$0.43 13.12$\pm$0.58 31.44$\pm$0.49 5.24$\pm$0.08 45.87$\pm$0.43 12.99$\pm$0.52
MEND 68.73$\pm$0.34 29.24$\pm$0.52 64.11$\pm$0.32 28.05$\pm$0.54 68.81$\pm$0.31 30.30$\pm$0.48 41.56$\pm$0.40 10.95$\pm$0.57 67.85$\pm$0.33 22.48$\pm$0.56
ROME 72.16$\pm$0.31 23.74$\pm$0.46 70.54$\pm$0.25 22.39$\pm$0.54 72.90$\pm$0.36 25.86$\pm$0.52 43.59$\pm$0.51 12.37$\pm$0.55 70.10$\pm$0.43 21.07$\pm$0.59
MEMIT 76.21$\pm$0.36 30.49$\pm$0.52 74.25$\pm$0.31 28.65$\pm$0.61 76.44$\pm$0.33 32.20$\pm$0.48 47.80$\pm$0.34 16.09$\pm$0.59 75.31$\pm$0.37 22.73$\pm$0.61
AlphaEdit 73.92$\pm$0.29 26.59$\pm$0.49 72.96$\pm$0.26 25.92$\pm$0.51 72.63$\pm$0.31 24.95$\pm$0.50 44.67$\pm$0.46 13.79$\pm$0.49 69.85$\pm$0.36 23.04$\pm$0.59
AnyEdit 97.76$\pm$0.11 92.96$\pm$0.24 96.60$\pm$0.19 95.60$\pm$0.35 97.76$\pm$0.14 95.87$\pm$0.23 62.63$\pm$0.44 46.51$\pm$0.59 96.33$\pm$0.21 94.32$\pm$0.23
UnKE 98.34$\pm$0.15 93.33$\pm$0.26 93.38$\pm$0.21 78.42$\pm$0.32 98.62$\pm$0.14 96.37$\pm$0.22 59.62$\pm$0.44 32.89$\pm$0.59 98.33$\pm$0.13 95.42$\pm$0.20
AnyEdit* 99.86$\pm$0.08 99.68$\pm$0.21 94.70$\pm$0.12 85.75$\pm$0.23 99.95$\pm$0.01 99.98$\pm$0.01 64.24$\pm$0.48 45.31$\pm$0.55 99.89$\pm$0.06 99.69$\pm$0.09
Qwen2.5-7B-It Pre-edited 64.18$\pm$0.37 25.88$\pm$0.59 64.39$\pm$0.34 24.02$\pm$0.55 65.50$\pm$0.34 18.24$\pm$0.60 44.74$\pm$0.41 17.29$\pm$0.51 67.71$\pm$0.37 19.58$\pm$0.49
FT-L 44.02$\pm$0.43 13.78$\pm$0.56 40.33$\pm$0.36 12.93$\pm$0.49 46.66$\pm$0.48 14.63$\pm$0.58 32.34$\pm$0.50 12.31$\pm$0.62 47.47$\pm$0.42 15.75$\pm$0.55
MEND 69.49$\pm$0.38 27.77$\pm$0.61 62.01$\pm$0.44 27.92$\pm$0.57 69.54$\pm$0.54 25.47$\pm$0.49 52.86$\pm$0.40 22.81$\pm$0.54 69.40$\pm$0.32 32.39$\pm$0.44
ROME 74.73$\pm$0.33 31.52$\pm$0.42 71.90$\pm$0.21 28.12$\pm$0.38 75.89$\pm$0.38 36.42$\pm$0.45 55.67$\pm$0.47 25.79$\pm$0.59 72.18$\pm$0.373 35.61$\pm$0.49
MEMIT 78.03$\pm$0.30 38.04$\pm$0.47 76.50$\pm$0.31 28.65$\pm$0.50 77.19$\pm$0.32 38.95$\pm$0.48 56.04$\pm$0.40 25.73$\pm$0.57 73.15$\pm$0.32 34.39$\pm$0.54
AlphaEdit 80.48$\pm$0.29 42.77$\pm$0.36 78.38$\pm$0.21 38.26$\pm$0.38 80.66$\pm$0.25 45.55$\pm$0.37 56.99$\pm$0.49 27.69$\pm$0.59 74.35$\pm$0.31 41.07$\pm$0.44
AnyEdit 98.05$\pm$0.16 94.89$\pm$0.29 93.56$\pm$0.15 79.98$\pm$0.28 98.08$\pm$0.15 95.09$\pm$0.19 65.40$\pm$0.38 43.49$\pm$0.47 98.14$\pm$0.13 96.39$\pm$0.18
UnKE 96.97$\pm$0.18 91.01$\pm$0.24 89.17$\pm$0.15 67.00$\pm$0.29 97.34$\pm$0.13 90.44$\pm$0.16 59.29$\pm$0.48 29.27$\pm$0.61 95.04$\pm$0.23 87.60$\pm$0.25
AnyEdit* 99.35$\pm$0.12 98.82$\pm$0.24 94.81$\pm$0.13 82.60$\pm$0.26 99.63$\pm$0.09 98.99$\pm$0.10 60.78$\pm$0.39 32.95$\pm$0.59 99.09$\pm$0.07 97.98$\pm$0.10
@@ 5.1 Experimental Setup
LLMs & Baseline Methods. We conducted experiments using two widely adopted LLMs: Llama3-8B-Instruct and Qwen2.5-7B-Instruct. For comparison with our method, we evaluated against several model editing methods, including FT-L (Zhu et al., 2020), MEND (Mitchell et al., 2022a), ROME (Meng et al., 2022), MEMIT (Meng et al., 2023), AlphaEdit (Fang et al., 2024), and UnKE (Deng et al., 2024).
Datasets and Evaluation Metrics. To evaluate the performance of unstructured long-form knowledge editing, we employed existing benchmarks, including UnKEBench (Deng et al., 2024) and AKEW (Wu et al., 2024). To further assess the editing performance across diverse-formatted knowledge, we constructed a dataset named EditEverything. For evaluation metrics, we assessed the similarity between the model’s edited outputs and the editing targets from three perspectives: original questions, paraphrased questions, and sub-questions. The evaluation was conducted using both semantic similarity (BERT Score (Zhang et al., 2019)) and lexical similarity (ROUGE Scores (Lin, 2004)) to determine the editing success rate.
@@ 6 Related Work
Model Editing for Knowledge Update. Traditional model editing methods modify a model’s knowledge by either altering a small subset of parameters or introducing external memory. These approaches enable knowledge updates, mitigate model hallucination (Huang et al., 2024a), or facilitate information injection (Chen et al., 2024). Parameter-Modifying Methods update existing model parameters to encode new knowledge. Meta-learning-based approaches, such as KE (Cao et al., 2021), MEND (Mitchell et al., 2022a) and InstructEdit (Zhang et al., 2024), train hypernetworks to predict parameter updates, with MEND improving efficiency via low-rank gradient decomposition. Locate-then-edit methods, including ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023), use causal tracing to identify knowledge-relevant parameters and update them via least-squares optimization. NSE (Jiang et al., 2024) mitigates catastrophic forgetting in model editing by updating only a subset of neuron parameters. AlphaEdit (Fang et al., 2024) extends this to lifelong editing with a null-space projection strategy. In contrast, Parameter-Preserving Methods introduce additional parameters or memory instead of modifying existing ones. ICE (Zheng et al., 2023) and DeCK (Bi et al., 2024) achieve parameter-free model editing through in-context learning, enabling knowledge updates without modifying the model’s parameters. SERAC (Mitchell et al., 2022b) retrieves updates from external memory, while T-Patcher (Huang et al., 2023) and CaliNet (Dong et al., 2022) allocate new neurons to encode knowledge. GRACE (Hartvigsen et al., 2023) replaces hidden states with discrete codebook values, and WISE (Wang et al., 2024) integrates parameterized memory for efficient knowledge merging.
Unstructured Knowledge Editing. Recent research focuses on editing unstructured knowledge with free text beyond structured triples. Wu et al. (2024) highlight the limitations of prior methods in evaluating unstructured text editing and introduce AKEW as a benchmark. UnKE (Deng et al., 2024) refines locate-then-edit methods by updating all parameters within a single layer, improving their capability to handle unstructured knowledge. They also introduce UnKEBench, a dedicated benchmark for unstructured knowledge editing. Huang et al. (2024b) propose a dynamic perception module to efficiently locate commonsense knowledge parameters, enabling free-text updates. However, while these methods handle unstructured and lengthy knowledge, they remain limited to factual knowledge. In contrast, our work extends editing capabilities to diverse-formatted knowledge, which encompasses various textual structures beyond factual statements.

### v3

@@ Abstract
Large language models (LLMs) often produce incorrect or outdated information, necessitating efficient and precise knowledge updates. Current model editing methods, however, struggle with long-form knowledge in diverse formats, such as poetry, code snippets, and mathematical derivations. These limitations arise from their reliance on editing a single token’s hidden state, a limitation we term as “efficacy barrier”. To solve this, we propose AnyEdit, a new autoregressive editing paradigm. It decomposes long-form knowledge into sequential chunks and iteratively edits the key token in each chunk, ensuring consistent and accurate outputs. Theoretically, we ground AnyEdit in the Chain Rule of Mutual Information, showing its ability to update any knowledge within LLMs. Empirically, it outperforms strong baselines by 21.5% on benchmarks including UnKEBench, AKEW, and our new EditEverything dataset for long-form diverse-formatted knowledge. Additionally, AnyEdit serves as a plug-and-play framework, enabling current editing methods to update knowledge with arbitrary length and format, significantly advancing the scope and practicality of LLM knowledge editing. Our code is available at: https://github.com/jianghoucheng/AnyEdit.
@@ 1 Introduction
Figure 1: Comparison of current methods and our AnyEdit. (a) and (d) illustrate the editing processes; (c) and (e) show the editing efficacy as the number of tokens within the to-be-updated knowledge increases; (b) and (f) depict the type of knowledge that each method can edit.
@@
Despite their success, existing model editing methods mostly face significant limitations in the length and diversity of the knowledge they can update. As shown in Figure 1 (c), even leading methods like AlphaEdit (Fang et al., 2025) and RECT (Gu et al., 2024) struggle to handle updates exceeding 100 tokens. Worse still, most methods are restricted to the knowledge represented as structured (subject, relation, object) triples. However, real-world knowledge is often encoded in diverse formats (e.g., mathematical derivations and code snippets as shown in Figure 1 (f)) and could exceed the 100-token threshold (Wu et al., 2024). These constraints are ill-suited for real-world scenarios, significantly narrowing the scope of model editing and hindering its broader advancement.
@@
Here we first conduct an in-depth analysis to identify why current methods fail for long-form diverse-formatted knowledge. Considering Figure 1 (a) again, existing methods typically rely on locating a single token, assuming that altering its hidden states will suffice to edit the LLM’s output (i.e., enabling the model to generate desired outputs reflecting new knowledge). However, long-form diverse-formatted knowledge is inherently more complex and information-dense than a single triplet, often requiring the integration of multiple critical tokens and intricate interdependencies among their hidden states. Thus, altering just a single token’s hidden state is insufficient to ensure consistent and accurate knowledge generation. We term this limitation as the efficacy barrier of single-token editing [^1] The “Efficacy” is the metric proposed by the model editing method ROME (Meng et al., 2022), aiming to evaluate the success rate of editing. For more details, please refer to Appendix A.2., which is empirically validated in Section 3. As a consequence, existing methods remain constrained by the paradigm of locating a single token and, as a result, are unable to overcome the aforementioned limitation.
@@
Hence, a critical question naturally arises for updating long-form diverse-formatted knowledge: “Can multiple tokens be jointly located and edited to enable complex knowledge updates?” A straightforward solution is to directly extend single-token editing to multiple tokens. However, it risks interference or conflicts between hidden state perturbations, undermining the coherence of to-be-updated knowledge and causing the performance drop as Figure 1 (c) showcases. In sight of this, we introduce AnyEdit, an autoregressive editing paradigm that enables collaborative multi-token updates. As illustrated in Figure 1 (d), AnyEdit decomposes long-form knowledge into sequential chunks, treating each as independent sub-knowledge. During editing, we iteratively (1) locate the final token of the current chunk and (2) perturb its hidden states to maximize the likelihood of generating the subsequent chunk. Building on the Chain Rule of mutual information (Dobrushin, 1963) in Information Theory (Kullback, 1997), we theoretically demonstrate that this autoregressive process ensures the generation of consistent, complete long-form knowledge. AnyEdit offers two key advantages: (1) Adaptivity: The number of edited tokens can be adaptively adjusted with the knowledge length, avoiding redundant edits. (2) Generality: It supports diverse knowledge formats (e.g., poetry, code, math) by decoupling structure-specific constraints. By shifting from single-token to multi-token collaborative editing, AnyEdit overcomes the challenge of efficacy barrier, demonstrating the potential for practical, updating complex knowledge of diverse formats.
@@
To validate AnyEdit, we conducted a comprehensive evaluation comparing it with leading model editing methods (e.g., MEMIT (Meng et al., 2023), AlphaEdit (Fang et al., 2025), and UnKE (Deng et al., 2025)) on the prevailing LLMs such as Llama3-8B-Instruct [^2] https://llama.meta.com/llama3 and Qwen2.5-7B-Chat (Yang et al., 2024). Beyond standard benchmark datasets (e.g., Counterfact (Meng et al., 2022) and ZsRE (Meng et al., 2023)) that represent knowledge as triples, we curate EditEverything, a new benchmark for long-form knowledge in diverse formats. As shown in Figure 1 (f), this dataset includes entries up to 458 tokens — over twice the length of the longest sequences in existing benchmarks (e.g., 156 tokens in AKEW (Wu et al., 2024)) — and spans multiple domains, including mathematics, news, code, and biochemistry. Results on EditEverything and standard benchmarks demonstrate that AnyEdit surpasses all baselines, achieving a 21.5% average improvement in editing accuracy with comparable computational overhead. Furthermore, as the first autoregressive editing framework, AnyEdit also enables seamless integration with existing locate-then-edit methods, as shown in Figure 1 (e). This plug-and-play capability equips traditional approaches with the ability to handle knowledge with arbitrary length and format, significantly broadening the scope and practicality of LLM knowledge editing.
@@
Figure 2: Relationship between knowledge format, original output probability, and efficacy when applying advanced editing methods to update triplet-structured and diverse-formatted knowledge. For each category, we randomly sample 200 knowledge instances to conduct experiments.
@@
Figure 3: Relationship between the number of tokens in to-be-updated knowledge, probability shift under random perturbations, and efficacy. We conduct experiments by truncating the sampled knowledge instances to enable editing across different token lengths. The lighter-colored bands represent variance.
@@ 2 Preliminary
where $\mathbb{P}_{f({\bm{h}}_{t}+\bm{\hat{\delta}})}$ represents the output probability when replacing ${\bm{h}}_{t}$ with $^{\bm{h}}_{t}+\bm{\hat{\delta}}$ in the LLM. Finally, the LLM parameters are updated such that, given the input $X$, the hidden state of the key token is aligned with ${\bm{h}}_{t}+\bm{\delta}$. For more details, please refer to Appendix B.1.
3 Limitations of Single-token Editing
Although single-token editing methods have been extensively studied in recent years (Fang et al., 2025; Wang et al., 2024; Deng et al., 2025; Gu et al., 2024; Wu et al., 2024), we argue that they suffer from updating long-form, diverse formatted knowledge.
3.1 Editing Diverse-formatted Knowledge
@@
3.2 Editing Long-form Knowledge
Recent studies suggest that although LLMs leverage attention mechanism (Wang & Komatsuzaki, 2021), the dependency between distant tokens weakens as their positional distance increases (Liu et al., 2024a). As a result, for long outputs $Y$ (e.g., $Y$ exceeding 100 tokens), perturbations applied to input tokens may have a diminishing influence on the later tokens within $Y$. In such cases, the shift of generation probability of $Y$ introduced by the perturbation $\bm{\delta}$ may be too small, making it insufficient to increase the probability of $Y$ above that of any other potential output.
To validate this, we repeat the experiments described in Section 3.1. Additionally, we use causal tracing (Meng et al., 2022), a common strategy in model editing, to quantify the shift in the generation probability of $Y$ introduced by the perturbation on $X$ (details of causal tracing can be found in Appendix B.1). Figure 3 illustrates the relationship between: the number of tokens in $Y$, the probability shift, and editing efficacy. The results demonstrate that long-form knowledge, which is less affected by single-token edits on the input, shows poor editing efficacy. In other words, the low probability shift introduced by single-token edits emerges as the inherent limitation in updating long-form knowledge.
@@ 3.2 Editing Long-form Knowledge
Here $\eta$ computes the average success rate of making $Y_{i}$ the most probable output for each input $X_{i}$ after editing.
@@ 5 Experiments
[…]
@@ 5.1 Experimental Setup
LLMs & Baseline Methods. We conducted experiments using two widely adopted LLMs: Llama3-8B-Instruct and Qwen2.5-7B-Instruct. For comparison with our method, we evaluated against several model editing methods, including FT-L (Zhu et al., 2020), MEND (Mitchell et al., 2022a), ROME (Meng et al., 2022), MEMIT (Meng et al., 2023), AlphaEdit (Fang et al., 2025), and UnKE (Deng et al., 2025).
Datasets and Evaluation Metrics. To evaluate the performance of unstructured long-form knowledge editing, we employed existing benchmarks, including UnKEBench (Deng et al., 2025) and AKEW (Wu et al., 2024). To further assess the editing performance across diverse-formatted knowledge, we constructed a dataset named EditEverything. For evaluation metrics, we assessed the similarity between the model’s edited outputs and the editing targets from three perspectives: original questions, paraphrased questions, and sub-questions. The evaluation was conducted using both semantic similarity (BERT Score (Zhang et al., 2019)) and lexical similarity (ROUGE Scores (Lin, 2004)) to determine the editing success rate.
@@ 6 Related Work
Model Editing for Knowledge Update. Traditional model editing methods modify a model’s knowledge by either altering a small subset of parameters or introducing external memory. These approaches enable knowledge updates, mitigate model hallucination (Huang et al., 2025), or facilitate information injection (Chen et al., 2024). Parameter-Modifying Methods update existing model parameters to encode new knowledge. Meta-learning-based approaches, such as MEND (Mitchell et al., 2022a), InstructEdit (Zhang et al., 2024) and RLedit (Li et al., 2025), train hypernetworks to predict parameter updates, with MEND improving efficiency via low-rank gradient decomposition. Locate-then-edit methods, including ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023), use causal tracing to identify knowledge-relevant parameters and update them via least-squares optimization. NSE (Jiang et al., 2024) mitigates catastrophic forgetting in model editing by updating only a subset of neuron parameters. AlphaEdit (Fang et al., 2025) extends this to lifelong editing with a null-space projection strategy. In contrast, Parameter-Preserving Methods introduce additional parameters or memory instead of modifying existing ones. ICE (Zheng et al., 2023) and DeCK (Bi et al., 2024) achieve parameter-free model editing through in-context learning, enabling knowledge updates without modifying the model’s parameters. SERAC (Mitchell et al., 2022b) retrieves updates from external memory, while T-Patcher (Huang et al., 2023) and CaliNet (Dong et al., 2022) allocate new neurons to encode knowledge. GRACE (Hartvigsen et al., 2023) replaces hidden states with discrete codebook values, and WISE (Wang et al., 2024) integrates parameterized memory for efficient knowledge merging.
Unstructured Knowledge Editing. Recent research focuses on editing unstructured knowledge with free text beyond structured triples. Wu et al. (2024) highlight the limitations of prior methods in evaluating unstructured text editing and introduce AKEW as a benchmark. UnKE (Deng et al., 2025) refines locate-then-edit methods by updating all parameters within a single layer, improving their capability to handle unstructured knowledge. They also introduce UnKEBench, a dedicated benchmark for unstructured knowledge editing. Huang et al. (2024) propose a dynamic perception module to efficiently locate commonsense knowledge parameters, enabling free-text updates. However, while these methods handle unstructured and lengthy knowledge, they remain limited to factual knowledge. In contrast, our work extends editing capabilities to diverse-formatted knowledge, which encompasses various textual structures beyond factual statements.